# Out-of-Distribution Detection by Leveraging Between-Layer Transformation Smoothness

### v1

@@ 4.5 Types of distribution shift
RoBERTa Semantic 61.61 69.46 69.50 69.41 68.34 69.36 68.91 70.56 72.03
Background 62.70 54.26 54.26 50.17 48.18 54.33 49.13 54.19 59.40
ELECTRA Semantic 62.49 63.17 63.12 60.92 62.14 63.23 65.67 62.45 64.22
Background 59.35 42.96 42.96 38.68 37.96 42.77 41.25 42.63 39.31
Table 4 shows the OOD detection performance on the semantic and background shift detection tasks. For the semantic shift, BLOOD exhibits suboptimal performance. However, in the case of the background shift, it notably outperforms all other methods, including the open-box approaches, some of which even perform worse than random. We suspect the subpar performance of other OOD detection methods in background shift detection may be attributed to models performing better on Yelp data compared to the sst data they were trained on (cf. Appendix C), because Yelp has longer texts with more semantic cues, making models more confident on OOD data. We speculate the discrepancy in performance between semantic and background shifts arises because BLOOD is focused on the encoding process of the query instances, while other methods only examine the model’s outputs. Consequently, BLOOD demonstrates greater sensitivity to the changes in the data-generating distribution. At the same time, other methods are better at detecting changes in the outputs, such as the introduction of an unknown class. In Appendix E we show that BLOOD is sensitive to the degree of distribution shift.
@@ 5 Conclusion
We have proposed a novel method for out-of-distribution (OOD) detection for deep neural networks (DNNs) called BLOOD. The method analyzes representation transformations across intermediate layers and requires only the access to model’s weights. Our evaluation on multiple text classification datasets using Transformer-based large pre-trained language models shows that BLOOD outperforms similar methods. Our analysis reveals that ID representations undergo smoother transformations between layers compared to OOD representations, because the model concentrates on ID region of the representation space during training. We demonstrated that the learning algorithm retains the original sharpness of the transformations of OOD intermediate representations for simpler datasets but increases the sharpness for more complex datasets. Future work includes applying BLOOD to other domains and developing a theoretical framework to explain the observed differences in between-layer smoothness between ID and OOD data.
@@ C.2 Datasets
Appendix D Dataset Complexity and Uncertainty
@@
Appendix E Degree of Distribution Shift
@@
Appendix F Additional Results

### v2

@@ 4.5 Types of distribution shift
RoBERTa Semantic 61.61 69.46 69.50 69.41 68.34 69.36 66.50 69.46 68.91 70.56 72.03
Background 62.70 54.26 54.26 50.17 48.18 54.33 50.46 49.32 49.13 54.19 59.40
ELECTRA Semantic 62.49 63.17 63.12 60.92 62.14 63.23 56.85 61.00 65.67 62.45 64.22
Background 59.35 42.96 42.96 38.68 37.96 42.77 40.66 38.53 41.25 42.63 39.31
Table 4 shows the OOD detection performance on the semantic and background shift detection tasks. For the semantic shift, BLOOD exhibits suboptimal performance. However, in the case of the background shift, it notably outperforms all other methods, including the open-box approaches, some of which even perform worse than random. We suspect the subpar performance of other OOD detection methods in background shift detection may be attributed to models performing better on Yelp data compared to the sst data they were trained on (cf. Appendix C), because Yelp has longer texts with more semantic cues, making models more confident on OOD data. We speculate the discrepancy in performance between semantic and background shifts arises because BLOOD is focused on the encoding process of the query instances, while other methods only examine the model’s outputs. Consequently, BLOOD demonstrates greater sensitivity to the changes in the data-generating distribution. At the same time, other methods are better at detecting changes in the outputs, such as the introduction of an unknown class. In Appendix G we show that BLOOD is sensitive to the degree of distribution shift.
@@ 5 Conclusion
We have proposed a novel method for out-of-distribution (OOD) detection for Transformer-based networks called BLOOD. The method analyzes representation transformations across intermediate layers and requires only the access to model’s weights. Our evaluation on multiple text classification datasets using Transformer-based large pre-trained language models shows that BLOOD outperforms similar methods. Our analysis reveals that ID representations undergo smoother transformations between layers compared to OOD representations, because the model concentrates on ID region of the representation space during training. We demonstrated that the learning algorithm retains the original sharpness of the transformations of OOD intermediate representations for simpler datasets but increases the sharpness for more complex datasets.
Acknowledgment
We thank the anonymous reviewers for their insightful comments. Our heartfelt appreciation goes to the members of TakeLab for their continuous support and valuable input. Special thanks to Nina Drobac and Stjepan Šebek for their feedback and helpful suggestions. This work has been supported by the Croatian Science Foundation under the project IP-2020-02-8671 PSYTXT (“Computational Models for Text-Based Personality Prediction and Analysis”).
@@ C.2 Datasets
Appendix D BLOOD for image classification
To explore the effectiveness of our BLOOD methodology on image data, we utilized CIFAR-10 and CIFAR-100 as our ID datasets (Krizhevsky et al., 2009). These were chosen for their resemblance to the ImageNet dataset, which was used to pre-train the Vision Transformer (ViT) (Dosovitskiy et al., 2021). We tested OOD detection using two different datasets: Street View House Numbers (SVHN; Yuval, 2011), which comprises real-world imagery of house numbers offering different visual features from CIFAR datasets, and the Beans dataset, which contains images relevant to agricultural disease classification, adding another layer of diversity to our evaluation.
In addition to the Transformer architecture, we also employ convolutional neural networks (CNNs), specifically ResNet34 and ResNet50 (He et al., 2016), to gauge how our method performs on other architectures. For CNNs, we employed a learning rate of $0.01$ and trained the models over $50$ epochs. In contrast, for our transformer-based model, the ViT, we adopted a fine-tuning approach, utilizing a learning rate of $2\times 10^{-5}$ with an additional weight decay of $0.001$, over a shorter span of $10$ epochs. This distinction in training methodologies is reflective of the different learning dynamics and capacities of CNNs versus Transformers.
[…]
Table 6: The performance of OOD detection methods on the task of OOD detection on image classification measured by AUROC (%). The best-performing method is in bold. Higher is better.
[…]
SVHN
ResNet34 CIFAR-10 84.00 49.99 88.92 89.71 88.23 81.39 82.29 83.15
CIFAR-100 87.05 49.51 73.64 76.20 77.90 78.31 85.97 80.54
ResNet50 CIFAR-10 84.13 49.91 89.55 90.29 84.36 90.07 88.01 88.20
CIFAR-100 79.77 50.50 79.86 83.32 82.70 83.42 81.80 80.89
ViT CIFAR-10 99.37 92.45 99.12 98.19 99.55 98.91 96.94 95.01
CIFAR-100 80.97 89.07 84.06 84.58 88.21 87.47 85.53 81.29
Beans
ResNet34 CIFAR-10 85.68 14.16 89.23 91.55 92.41 91.09 81.04 90.67
CIFAR-100 86.40 23.54 73.32 79.90 81.22 83.97 76.55 81.02
ResNet50 CIFAR-10 78.37 50.28 79.18 80.45 79.02 79.62 69.33 77.26
CIFAR-100 86.62 49.76 79.13 80.79 82.31 83.07 85.29 75.16
ViT CIFAR-10 95.41 99.58 99.31 99.31 99.98 98.79 96.92 94.31
CIFAR-100 91.68 96.53 96.02 95.71 95.81 95.83 89.33 88.14
Appendix E BLOOD as an Open-Box Method
In the scenarios where more resources are available, beyond just the trained model, it is possible to utilize those resources to improve the OOD detection performance of BLOOD. Similarly to a lot of popular OOD detection methods (Liang et al., 2018; Sun et al., 2021; Sun & Li, 2022), BLOOD can be improved by using a small validation set to learn the optimal weights for the weighted average of the BLOOD score in each layer.
[…]
Model SST SUBJ AGN TREC BP AR MG NG
RoBERTa 73.94 82.98 81.39 91.73 93.47 96.25 92.46 89.97
ELECTRA 77.67 77.30 82.76 98.73 96.54 91.97 90.92 85.19
Appendix F Dataset Complexity and Uncertainty
@@
Appendix G Degree of Distribution Shift
@@
Appendix H Additional Results